# When LLMs Meet Cunning Questions: A Fallacy Understanding Benchmark for Large Language Models

### v1

@@ Abstract
Recently, Large Language Models (LLMs) have made remarkable evolutions in language understanding and generation. Following this, various benchmarks for measuring all kinds of capabilities of LLMs have sprung up. In this paper, we challenge the reasoning and understanding abilities of LLMs by proposing a FaLlacy Understanding Benchmark (FLUB [^1] Our data and codes are available at https://github.com/THUKElab/FLUB containing cunning questions that are easy for humans to understand but difficult for models to grasp. Specifically, the cunning questions that FLUB focuses on mainly consist of the tricky, humorous, and misleading questions collected from the real internet environment. And we design three tasks with increasing difficulty in the FLUB benchmark to evaluate the fallacy understanding ability of LLMs. Based on FLUB, we investigate the performance of multiple representative and advanced LLMs, reflecting our FLUB is challenging and worthy of more future study. Interesting discoveries and valuable insights are achieved in our extensive experiments and detailed analyses. We hope that our benchmark can encourage the community to improve LLMs’ ability to understand fallacies.
When LLMs Meet Cunning Questions:
A Fallacy Understanding Benchmark for Large Language Models
[…]
@@ 1 Introduction
Large Language Models (LLMs) have shown great abilities to understand human languages, including information extraction Yu et al. (2023), text correction Li et al. (2023b), complex reasoning Bang et al. (2023), etc. Researchers have constructed numerous benchmarks to test the capabilities of LLMs in various aspects Ma et al. (2022); Huang et al. (2022, 2023); Li et al. (2023c, a). By using collected questions to interact with LLMs, researchers can analyze the behavior of LLMs to compare the performance of different LLMs and study how to further improve LLMs.
@@
Although many LLM benchmarks have sprung up, we believe that existing benchmarks are not challenging enough to truly measure the human-like intelligence of LLMs. In particular, we are still wondering whether LLMs can understand cunning questions that may contain misleading, wrong premise, intentional ambiguity, and so forth, considering that almost all LLMs are trained on “cleaned” and “correct” corpora. Therefore, we build a FaLlacy Understanding Benchmark (FLUB) to challenge LLMs for solving these problems.
@@
Figure 1: The examples of how LLMs and humans perform when faced with cunning questions. The LLM we use is ChatGPT-3.5 on Jan 23, 2024.
Figure 2: The data annotation example of FLUB.
@@
Figure 1 shows the running examples from FLUB. From these cases, we directly feel the different behaviors of LLMs and humans when facing cunning questions. In the first example, LLMs ignore the common sense that the lotus root itself has many holes in its structure and fall into the trap of the cunning question, wrongly judging that the holes in the lotus root are caused by insect infestation. In the second example, LLMs fail to see the logic that depositing money into random ATMs does not create problems and therefore give an answer that seems reasonable but is absurdly laughable. In fact, these cunning questions for LLMs are very easy to handle for human intelligence. Therefore, it is very urgent and meaningful to construct a benchmark composed of cunning questions to evaluate and thereby promote the improvement of LLMs’ fallacy understanding capabilities.
@@
Inspired by the above motivation, we collect real cunning questions as our raw data from a famous Chinese online forum, the “Ruozhi Bar” (retard forum)  [^2] https://tieba.baidu.com/f?kw=%E5%BC%B1%E6%99%BA&ie=utf-8. This forum is popular for its cunning and unreasonable posts, which are generally easy for humans to understand but challenging for LLMs. The characteristics of the posts contained in this forum are consistent with our research motivation, so choosing it as the data source well supports FLUB’s evaluation of LLMs’ fallacy understanding ability. After data cleaning and annotating of question types, FLUB has 8 fine-grained types of cunning questions and most of the questions in FLUB fall into two types of fallacy, namely, faulty reasoning and word game. Moreover, we also manually annotated one correct answer (i.e., the explanation of the question) and three confusing wrong answers for each question in FLUB, as shown in Figure 2.
@@
Based on our constructed FLUB and its annotation information, we design three tasks with increasing difficulty to test whether the LLMs can understand the fallacy text and solve the “cunning” questions. Specifically, (1) Answer Selection: The model is asked to select the correct one from the four answers provided by FLUB for each input question. (2) Question Type Classification: Given a cunning question as input, the model is expected to directly identify its fallacy type defined in our scheme. (3) Question Explanation: We hope the model sees a cunning question and intelligently generates a correct explanation for the question, just like humans, without falling into its trap.
@@
In our experiments, we select representative and advanced LLMs to be evaluated on FLUB. Our empirical study reveals: (1) LLMs are very poor in their ability to perceive fallacy types in cunning questions. (2) For a specific task, LLMs with larger parameter sizes do not always perform better. (3) There is a close relationship between the Answer Selection task and the Question Explanation task, and the interaction between them is critical to promoting the understanding of fallacies in LLMs. (4) On FLUB, the widely used Chain-of-Thought and In-context Learning techniques deserve further improvement and research. We believe that our proposed FLUB and all our findings are crucial for LLMs to comprehend the fallacy and handle cunning questions in the real world.
Question Type # of Samples Example
错误类比 11 很多人出门后担心刚刚没有关门,为什么进门后不担心刚刚没有开门?
False Analogy Many people worry about forgetting to close the door when they leave home.
Why don’t they worry about whether they have opened the door when they come in?
冷笑话 44 忘记把钱存在哪个ATM机里了怎么办?银行好几台ATM机,还长得都一样。
Lame Jokes What should I do if I forget which ATM I deposited money into?
The bank has several ATMs, and they all look the same.
字音错误 5 因为美国队长,小明每次在美国排队都要排一个多小时。
Phonetic Error Because of Captain America (also read as “long queues in America” in Chinese),
Xiao Ming has to wait over an hour whenever he queues in the U.S.
歧义 35 语文老师说我写的句子是病句,我应该给这个病句吃头孢,还是打点滴呢?
Ambiguity My teacher said the sentence is grammatically incorrect (“sick sentence” in Chinese).
Should I give this sentence some antibiotics or administer an IV drip?
悖论 29 “凡事无绝对”这句话过于绝对。
Paradox The phrase “Nothing is absolute” is too absolute.
事实性错误 12 一吨的铁和一吨的棉花哪个重啊?
Factual Error Which one weighs more, a ton of iron or a ton of cotton?
推理错误 445 根据我在养老院的调查数据,我国的人口老龄化已经相当严重了。
Reasoning Error According to my survey data from nursing homes,
the aging of the population in our country has become quite severe.
文字游戏 239 人类70%是水,所以10个人里有7个人是水伪装成的人!
Word Game 70% of the human body is water, so 7 out of 10 people are water disguised as humans!
未分类 24 在高速路的服务区开酒吧有可行性吗?
Undefined Is it feasible to open a bar at a highway service area?
Table 1: Question types of FLUB and corresponding examples.
@@ 2.1 Benchmark Construction
We collect raw cunning question data from “Ruozhi Bar” in Baidu Tieba  [^3] https://tieba.baidu.com. “Ruozhi Bar” is one of the most famous online forums in the Chinese internet community, and people often post some interesting or “silly” questions on it just for fun. We find that many of the posts on this forum are tricky questions or brain-teaser-like texts, which is exactly in line with our purpose of using cunning questions to challenge LLMs, so we decide to utilize this forum as our data source. As a result of automatic crawling, we initially collect 9,927 candidate posts, including the title, body text, and the first comment of the post. Notably, according to the Baidu Bar agreement  [^4] https://baike.baidu.com/item/%E8%B4%B4%E5%90%A7%E5%8D%8F%E8%AE%AE/8397765, the data on Baidu Tieba can be used for academic research free of charge and without liability.
@@
We employ annotators to manually filter out irrelevant posts that do not present tricky or cunning questions. Since the collected original posts contain irrelevant content such as links and pictures, we also require data annotators to extract the fallacious and illogical contents from the title and body text of each post and rewrite them into a complete question. Besides, it is worth noting that we carefully ensure that the questions remained in FLUB are ethical texts. This process includes user information anonymization, sensitive information removal, and filtering of impolite posts. In total, we obtain 844 data samples to form FLUB.
@@
To ensure the annotation quality of FLUB and taking into account the characteristics of our study, when we select annotators, our criteria for selecting annotators is that the person must be a native Chinese speaker and have a bachelor’s degree. The detailed annotation workflows for each type of information are as follows:
@@
    Question Type Annotation: To do this kind of annotation, we first define 8 question types within the collected questions along with their corresponding examples. Subsequently, each data sample is processed by three junior annotators, who are required to select an appropriate question type for the sample. We achieve the initial annotation results based on the voting results among three annotators. The initial annotation results become the final annotation information after being reviewed by the senior annotator (and modified if necessary). The schema for the types is shown in Table 1.
@@
    Correct Explanation Annotation: We assign two junior annotators to write the explanation or answer for each sample independently. We ask them to try to explain the given question in a detailed, objective, and unambiguous way. The senior annotator then selects (and modifies if necessary) the more suitable text written by the two junior annotators.
@@
    Wrong Candidates Annotation: This part annotation is to obtain the wrong candidate answer that may be likely to be answered incorrectly for each question. We assign three junior annotators for each sample and require each of them to write three different incorrect answers based on their understanding of the question. Particularly, we emphasize to each junior annotator that the three different wrong answers they write should ensure diversity and resemble as much as possible the answers that LLMs can easily produce. For each sample’s nine initial incorrect answers, the senior annotator selects the three most challenging sentences as the final wrong candidates.
It is worth mentioning that we have prepared sufficient and representative samples for annotators to learn and pre-annotate to ensure that they fully understand the information we want to annotate before they officially start annotation. Our entire annotation process lasted 2 weeks. Other annotation details are presented in Appendix A
@@ 2.2 Dataset Analysis
As shown in Table 1, most data in FLUB belong to the types of reasoning errors and word games. This is because these two types of questions appear widely in “Ruozhi Bar” forum whose purpose is to challenge human intelligence. A large number of cunning questions involving reasoning errors and word games ensure that FLUB is challenging enough. Besides, we observe that some types of questions are relatively rare, such as phonetic errors. In fact, this is because our data come entirely from the real world and are all carefully constructed by netizens. Cases of cunning questions caused by phonetic errors are indeed rare in the real world. From another perspective, the data distribution also reflects that FLUB is real and close to human lives, so it can better measure the intelligence gap between humans and LLMs.
@@
Since question-type annotation is essentially a classification process performed by multiple annotators, we analyze the annotation quality of this information. Specifically, we calculate Fleiss’ Kappa Falotico and Quatto (2015) to reflect the three junior annotator’s Inter-Annotator Agreement (IAA). Our final obtained Fleiss Kappa result is greater than 0.767, which shows that our annotation results have excellent consistency and quality Landis and Koch (1977).
@@ 2.3 Benchmark Task Setups
To evaluate the fallacy understanding ability of LLMs, we design three benchmark tasks on FLUB: Answer Selection, Question Type Classification, and Question Explanation. For each task, we design prompts to guide LLMs on the expected output. Particularly, for Task 1 and Task 2, to stimulate the reasoning ability of LLMs, we design prompts with the Chain-of-Thought idea Wei et al. (2022) as shown in Figure 3. For Task 3, we believe that the task goal itself is straightforward enough, so it is not suitable for the Chain-of-Thought. Below we describe the details of the three benchmark tasks:
Figure 3: Our designed prompts for FLUB. Task 3(a) is for the questions that are not expressed in the form of inquiries. Task 3(b) is for inquiries. Note that here we show Chain-of-Thought prompts for Task 1 and Task 2, and their prompts without Chain-of-Thought are in Appendix B. The English translations of our prompts are also in Appendix B.
@@
In Task 1, LLMs are required to select the correct answer from four given candidate explanations for each question. The annotation of candidate explanations is illustrated in Figure 2. In general, each sample in this task is a tuple $\{p,q,O_{A},O_{B},O_{C},O_{D},l\}$, where $p$ is our given prompt as shown in Figure 3, $q$ is the input question, $O_{A}$, $O_{B}$, $O_{C}$, and $O_{D}$ are four candidate explanations, and $l\in\{A,B,C,D\}$ is the golden label indicating $O_{l}$ is the correct explanation. The design motivation of this task is to test whether LLMs can distinguish right from wrong when seeing the correct and wrong answers in the context of a given cunning question.
Task 2: Question Type Classification
@@
If LLMs are directly tasked with determining the corresponding question type, it will help us in conducting an initial automated assessment of the LLM’s understanding ability. The question type classification task is specifically designed to evaluate whether LLMs can classify the cunning question into categories aligned with human intuition based on the hidden irrational aspects within the current question. The annotated problem types are shown in Table 1. During task evaluation, all the problem types will be combined with the prompt to allow LLMs to directly pick the correct type of cunning question. We believe that LLMs understanding the concept or type of fallacy first is an unavoidable prerequisite for them to handle cunning questions well.
Task 3: Question Explanation
@@
To further test whether LLMs truly understand the given question, we design the explanation task. In this task, the designed prompt and questions are directly input into LLMs, enabling them to “read” input questions and generate corresponding explanations. Note that since some of the questions are not expressed in the form of inquiries, we have additionally set a prompt to guide LLMs in identifying the question (See Figure 3). The generated explanations will be compared with the correct explanation for evaluation. If LLMs can generate correct explanations, we believe that they have the ability to identify the traps of cunning questions and have come close to human intelligence.
Evaluation Metrics
For Task 1 and Task 2, we automatically calculate accuracy directly based on the LLMs’ selection and classification.
To evaluate the quality of LLMs’ generated explanations in Task 3, we employ automated evaluation along with human evaluation to score their explanations respectively. For automated evaluation, inspired by MT-Bench Zheng et al. (2023), we construct prompts that incorporate the task instruction, questions, LLM’s explanations and reference answers. These prompts are fed into GPT-4, which is tasked with assigning a score ranging from 1 to 10. The prompt for the automated evaluation is illustrated in Appendix C. For human evaluation, we hire 3 evaluation annotators to rate LLMs’ explanations, with scores ranging from $\{1,2,3,4,5\}$. To ensure fair evaluation of the explanations of LLMs, we developed a set of scoring guidelines for annotators, including the definitions and relevant examples for each score. The scoring guidelines of human evaluation are presented in Appendix D.
Models Open Selection Classification Explanation
Source Accuracy Accuracy GPT-4 Score
ERNIE-Bot-3.5 Baidu (2023) 52.76 (38.37) 21.71 (16.59) 6.349
ERNIE-Bot-3.5-Turbo Baidu (2023) 32.97 (34.65) 1.71 (10.12) 5.783
ERNIE-Bot-4.0 Baidu (2023) 75.66 (71.34) 20.00 (12.32) 7.729
GPT-3.5-Turbo OpenAI (2023) 50.48 (48.08) 5.61 (7.68) 6.233
GPT-4-Turbo OpenAI (2023) 79.38 (82.73) 15.37 (15.00) 8.952
ChatGLM3-6B Du et al. (2022) 35.01 (48.44) 17.56 (18.54) 4.983
Qwen-7B-Chat Bai et al. (2023) 38.49 (34.17) 19.27 (24.88) 5.392
Qwen-14B-Chat Bai et al. (2023) 42.57 (39.69) 17.68 (18.78) 5.241
Qwen-72B-Chat Bai et al. (2023) 58.63 (59.35) 15.12 (15.49) 7.335
Yi-6B-Chat 01-AI (2023) 32.61 (36.57) 12.80 (17.80) 5.731
Yi-34B-Chat 01-AI (2023) 47.96 (61.15) 7.20 (20.73) 6.970
Baichuan2-7B-Chat Yang et al. (2023) 43.17 (36.45) 2.44 (6.34) 5.476
Baichuan2-13B-Chat Yang et al. (2023) 37.05 (40.41) 4.02 (4.15) 5.787
Random - 25.00 $1$
Table 2: The main results on FLUB. The results in parentheses are the performance with Chain-of-Thought. We bold the optimal and underline the suboptimal of closed-source and open-source models for convenience.

### v2

# When LLMs Meet Cunning Texts: A Fallacy Understanding Benchmark for Large Language Models

@@ Abstract
Recently, Large Language Models (LLMs) make remarkable evolutions in language understanding and generation. Following this, various benchmarks for measuring all kinds of capabilities of LLMs have sprung up. In this paper, we challenge the reasoning and understanding abilities of LLMs by proposing a FaLlacy Understanding Benchmark (FLUB) containing cunning texts that are easy for humans to understand but difficult for models to grasp. Specifically, the cunning texts that FLUB focuses on mainly consist of the tricky, humorous, and misleading texts collected from the real internet environment. And we design three tasks with increasing difficulty in the FLUB benchmark to evaluate the fallacy understanding ability of LLMs. Based on FLUB, we investigate the performance of multiple representative and advanced LLMs, reflecting our FLUB is challenging and worthy of more future study. Interesting discoveries and valuable insights are achieved in our extensive experiments and detailed analyses. We hope that our benchmark can encourage the community to improve LLMs’ ability to understand fallacies. Our data and codes are available at https://github.com/THUKElab/FLUB.
@@ 1 Introduction
Large Language Models (LLMs) have shown great abilities to understand human languages, including information extraction [1], text correction [2], humor understanding [3], etc. Researchers have constructed numerous benchmarks to evaluate LLMs in various aspects [4, 5, 6, 7, 8]. By using constructed benchmarks to interact with LLMs, researchers can analyze the behavior of LLMs to compare the performance of different LLMs and study how to further improve LLMs in a targeted manner.
@@
Although many LLM benchmarks have sprung up, we believe that existing benchmarks are not challenging enough to truly measure the human-like intelligence of LLMs. In particular, we are still wondering whether LLMs can understand cunning texts that may contain misleading, wrong premise, intentional ambiguity, and so forth, considering that almost all LLMs are trained on “cleaned” and “correct” corpora. Therefore, we build a FaLlacy Understanding Benchmark (FLUB) to challenge LLMs for solving these problems.
@@
(a) The examples of how LLMs and humans perform when faced with cunning texts. The LLM we use is ChatGPT-3.5 on Jan 23, 2024.
(b) We design three tasks, namely Cunning Type Classification, Fallacy Explanation, and Answer Selection (i.e., Multiple Choice).
Figure 1: The running examples and annotation examples of FLUB.
@@
Figure 1(a) shows the running examples from FLUB. From these cases, we directly feel the different behaviors of LLMs and humans when facing cunning texts. In the first example, LLMs ignore the common sense that the lotus root itself has many holes in its structure and fall into the trap of the cunning text, wrongly judging that the holes in the lotus root are caused by insect infestation. In the second example, LLMs fail to see the logic that depositing money into random ATMs does not create problems and therefore give an answer that seems reasonable but is absurdly laughable. In fact, these cunning texts for LLMs are very easy to handle for human intelligence. Therefore, it is very urgent and meaningful to construct a benchmark composed of cunning texts to evaluate and thereby promote the improvement of LLMs’ fallacy understanding capabilities.
@@
Inspired by the above motivation, we collect real cunning texts as our raw data from a famous Chinese online forum, the “Ruozhiba” (retard forum)  [^1] https://tieba.baidu.com/f?kw=%E5%BC%B1%E6%99%BA&ie=utf-8. This forum is popular for its cunning and unreasonable posts, which are generally easy for humans to understand but challenging for LLMs. The characteristics of the posts contained in this forum are consistent with our research motivation, so choosing it as the data source well supports FLUB’s evaluation of LLMs’ fallacy understanding ability. After data cleaning and annotating of cunning types, FLUB has 8 fine-grained types of cunning texts and most of the texts in FLUB fall into two types of fallacy, namely, faulty reasoning and word game. Moreover, we also manually annotated one correct answer (i.e., the explanation of the cunning text) and three confusing wrong answers for each input text in FLUB, as shown in Figure 1(b).
@@
Based on our constructed FLUB and its annotation information, we design three tasks with increasing difficulty to test whether the LLMs can understand the fallacy and solve the “cunning” texts. Specifically, (1) Answer Selection: The model is asked to select the correct one from the four answers provided by FLUB for each input text. (2) Cunning Type Classification: Given a cunning text as input, the model is expected to directly identify its fallacy type defined in our scheme. (3) Fallacy Explanation: We hope the model sees a cunning text and intelligently generates a correct explanation for the fallacy contained in the text, just like humans, without falling into its trap.
@@
In our experiments, we select representative and advanced LLMs to be evaluated on FLUB. Our empirical study reveals: (1) LLMs are very poor in their ability to perceive fallacy types in cunning texts. (2) For a specific task, LLMs with larger parameter sizes do not always perform better. (3) There is a close relationship between the Answer Selection task and the Fallacy Explanation task, and the interaction between them is critical to promoting the understanding of fallacies in LLMs. (4) On FLUB, the widely used Chain-of-Thought and In-context Learning techniques deserve further improvement and research. We believe that our proposed FLUB and all our findings are crucial for LLMs to comprehend the fallacy and handle cunning texts in the real world.
Figure 2: The definitions and examples of the cunning types in FLUB.
@@ 2.1 Benchmark Construction
We collect raw text data from “Ruozhiba” in Baidu Tieba  [^2] https://tieba.baidu.com. “Ruozhiba” is one of the most famous online forums in the Chinese internet community, and people often post interesting or “silly” texts on it just for fun. In addition, the recent study [9] also shows that the Ruozhiba data is very useful for improving the ability of Chinese LLMs. We find that many of the posts on this forum are tricky texts or brain-teaser-like texts, which is exactly in line with our purpose of using cunning texts to challenge LLMs, so we utilize this forum as our data source. As a result of automatic crawling, we initially collect 9,927 candidate posts. Notably, according to the Baidu Bar agreement  [^3] https://baike.baidu.com/item/%E8%B4%B4%E5%90%A7%E5%8D%8F%E8%AE%AE/8397765, the data on Baidu Tieba can be used for academic research free of charge and without liability.
@@
We employ annotators to manually filter out irrelevant posts that do not present cunning texts. Since the collected original posts contain irrelevant content such as links and images, we also require annotators to extract the fallacious and illogical contents from the raw post and rewrite them into a complete sentence. Besides, it is worth noting that we carefully ensure that the texts in FLUB are ethical texts. This process includes user information anonymization, sensitive information removal, and filtering of impolite posts. In total, we obtain 834 data samples to form FLUB.
@@
To ensure the annotation quality, our criteria for selecting annotators is that the person must be a native Chinese speaker and have a bachelor’s degree. In addition, because FLUB comes from the online forum, we also require annotators to have more than five years of experience as netizens. The detailed annotation workflows are as follows:
@@
    Cunning Type Annotation: We first define 8 cunning types within the collected texts along with their corresponding examples, as shown in Figure 2. Subsequently, each data sample is processed by three junior annotators, who are required to select an appropriate cunning type for the sample. We achieve the initial annotation results based on the voting results among three annotators. The initial annotation results become the final annotation information after being reviewed by the senior annotator (and modified if necessary).
@@
    Correct Explanation Annotation: We assign two junior annotators to write the explanation or answer for each sample independently. We ask them to try to explain the given text in a detailed, objective, and unambiguous way. The senior annotator then selects (and modifies if necessary) the more suitable text written by the two junior annotators.
@@
    Wrong Candidates Annotation: This part annotation is to obtain the wrong candidate answer that may be likely to be answered incorrectly for each input text. We assign three junior annotators for each sample and require each of them to write three different incorrect answers based on their understanding of the text. Particularly, we emphasize to each junior annotator that the three different wrong answers they write should ensure diversity and resemble as much as possible the answers that LLMs can easily produce. For each sample’s nine initial incorrect answers, the senior annotator selects the three most challenging sentences as the final wrong candidates.
Since the annotation difficulty of different information is different, the salary we pay to the annotators we employ is also different. Specifically, we pay each person who annotates the cunning type \$0.5 per sample, each person who writes the correct explanation \$1 per sample, and each person who writes the wrong candidates \$2 per sample. In addition to the junior annotators providing the initial annotation results, we also set three senior annotators with a salary of \$2 per sample, who are responsible for carefully checking the correctness of the annotation results provided by the junior annotators.
It is worth mentioning that we have prepared sufficient and representative samples for annotators to learn and pre-annotate to ensure that they fully understand the information we want to annotate before they officially start annotation. Our entire annotation process lasted 2 weeks.
@@ 2.2 Dataset Analysis
Data Size
FLUB comprises 834 samples that span 8 cunning types. It is worth emphasizing that the data size is not directly related to the evaluation effectiveness of a LLM benchmark. For example, TruthfulQA [10] and FreshQA [11], these benchmarks that have been widely used and had deep impacts, only have 817 and 500 test samples respectively. The main reasons limiting the size of FLUB are that it is derived entirely from real-world online forum posts and our rigorous high-quality data cleaning process, which retained 834 final samples from 9,927 candidate posts.
@@
As for the cunning type distribution of FLUB, most data in FLUB belong to the types of reasoning errors (53.4%) and word games (28.7%). This is because these two types of posts appear widely in “Ruozhi Bar” forum whose purpose is to challenge human intelligence. A large number of cunning texts involving reasoning errors and word games ensure that FLUB is challenging enough. Besides, we observe that some types of texts are relatively rare, such as phonetic errors (0.6%). In fact, this is because our data come entirely from the real world and are all carefully constructed by netizens. Cases of cunning texts caused by phonetic errors are indeed rare in the real world. To eliminate the impact of type imbalance when FLUB evaluating LLMs, we choose the F-1 score as the evaluation metric which comprehensively considers the type coverage.
@@
Since cunning type annotation is essentially a classification process performed by multiple annotators, we analyze the annotation quality of this information. Specifically, we calculate Fleiss’ Kappa [12] to reflect the three junior annotator’s Inter-Annotator Agreement (IAA). Our final obtained Fleiss Kappa result is greater than 0.767, which shows that our annotation results have excellent consistency and quality [13].
@@ 2.3 Benchmark Task Setups
To evaluate the fallacy understanding ability of LLMs, we design three benchmark tasks on FLUB: Answer Selection, Cunning Type Classification, and Fallacy Explanation. For each task, we design prompts to guide LLMs on the expected output. We also explore the prompting strategies of Chain-of-Thought and In-context Learning to conduct in-depth exploration on FLUB. The details of our designed prompts are shown in Appendix A. Below we introduce the details of our three tasks:
@@
In Task 1, LLMs are required to select the correct answer from four given candidate explanations for each input text. The annotation of candidate explanations is illustrated in Figure 1(b). In general, each sample in this task is a tuple $\{p,q,O_{A},O_{B},O_{C},O_{D},l\}$, where $p$ is our given prompt as shown in Appendix A, $q$ is the input text, $O_{A}$, $O_{B}$, $O_{C}$, and $O_{D}$ are four candidate explanations, and $l\in\{A,B,C,D\}$ is the golden label indicating $O_{l}$ is the correct explanation. The design motivation of this task is to test whether LLMs can distinguish right from wrong when seeing the correct and wrong answers in the context of a given cunning text.
Task 2: Cunning Type Classification
@@
If LLMs are directly tasked with determining the corresponding cunning type, it will help us in conducting an initial automated assessment of the LLM’s understanding ability. The cunning type classification task is specifically designed to evaluate whether LLMs can classify the cunning text into categories aligned with human intuition based on the hidden irrational aspects within the current text. The annotated problem types are shown in Figure 2. During task evaluation, all the problem types will be combined with the prompt to allow LLMs to directly pick the correct type of cunning text.
Task 3: Fallacy Explanation
@@
To further test whether LLMs truly understand the given cunning text, we design the explanation task. In this task, the designed prompt and input texts are directly input into LLMs, enabling them to “read” input texts and generate corresponding explanations. Note that since some texts are not expressed in the form of inquiries, we also set a prompt to guide LLMs in identifying the question (See Appendix A). The generated explanations will be compared with the correct explanation for evaluation. If LLMs can generate reasonable explanations, we believe that they have at least developed the ability to identify and avoid the traps of cunning texts.
Automatic Evaluation Metrics
For Task 1, we calculate Accuracy directly based on the LLMs’ selection results. For Task 2, considering that there are a few cunning types in FLUB with small sample size, we choose the F-1 Score to measure the performance of LLMs because it focuses on both the accuracy of model prediction and the coverage for positive class samples, thereby effectively avoiding bias caused by type imbalance and ensuring the rationality and reliability of evaluation. To evaluate the quality of LLMs’ generated explanations in Task 3, inspired by MT-Bench [14], we construct prompts that incorporate the task instruction, input texts, LLM’s explanations, and reference answers. These prompts are fed into GPT-4, which is tasked with assigning a GPT-4 Score ranging from 1 to 10. The prompt for the automated evaluation is illustrated in Appendix B.
Human Evaluation Settings
For Task 1 and Task 2, we conduct human evaluations to explore how well human-level intelligence could perform these two tasks. To ensure the fairness of the comparison between humans and LLMs, we hire 3 new persons who do not participate in the construction process of FLUB. After briefly introducing them to the objectives of Task 1 and Task 2 (without introducing additional knowledge and information), let them directly carry out selection and classification. For the human evaluation of Task 3, we mainly want to verify the effectiveness of the automatic GPT-4 score we use, therefore, we hire 3 evaluation annotators to rate LLMs’ explanations, with scores ranging from $\{1,2,3,4,5\}$. To ensure an accurate evaluation of the explanations of LLMs, we developed a set of scoring guidelines for annotators, including the definitions and relevant examples for each score. The scoring guidelines of human evaluation are presented in Appendix C.